# Quasinormal modes of spontaneous scalarized Kerr black holes

### v1

@@ Abstract
Recent studies reveal that Kerr black holes can undergo spontaneous scalarization via an appropriate coupling between a scalar field and the Gauss-Bonnet term. This mechanism introduces observable signatures in the quasinormal mode spectrum, which characterizes the ringdown phase of gravitational waves. In this work, we present, for the first time, the quasinormal modes of rotating scalarized black holes within the framework of Einstein-scalar-Gauss-Bonnet theory, without resorting to perturbative expansion of either the coupling constant or the black hole spin. Using spectral decomposition, we numerically construct both the background solution and the perturbation functions, reformulating the calculation of quasinormal modes as a generalized eigenvalue problem. Our results pave the way for advanced ringdown tests of spontaneous scalarization using gravitational wave data.
@@
Introduction — The emerging field of gravitational wave (GW) astronomy offers an unprecedented opportunity to explore fundamental physics in strong gravity regime LIGOScientific:2018mvr . As GW data grows and detector sensitivity improves, General Relativity (GR) – the most successful theory of gravity – can be tested from various perspectives Berti:2015itd . One of GR’s most remarkable predictions, black holes (BHs), plays a key role in probing strong gravity due to their simplicity Barack:2018yly . According to the no-hair theorem Herdeiro:2015waa , BHs in GR are uniquely characterized by mass, spin, and electric charge. Current and future GW observations enable precise tests of this theorem, including the potential existence of extra scalar fields or other matter fields that hint at new physics Isi:2019aib ; Barausse:2020rsu ; Brito:2017wnc ; Hannuksela:2018izj .
Spontaneous scalarization is an intriguing mechanism that introduces an extra scalar degree of freedom for BHs through the coupling of a scalar field with a source term, such as curvature or matter Antoniou:2017acq ; Doneva:2017bvd ; Antoniou:2017hxj ; Doneva:2018rou ; Collodel:2019kkx ; Herdeiro:2020wei ; Silva:2017uqg ; Berti:2020kgk ; Herdeiro:2018wub ; Silva:2017uqg ; Cunha:2019dwb . The spontaneous scalarization model can withstand rigorous tests in weak gravity regime wherein this model is indistinguishable from GR. On the other hand, this model allows a sharp transition to a stationary scalarized BH that can differ significantly from GR in the strong field regime, and hence the GWs emitted by the scalarized BHs may possess detectable deviations from GR expectations. For recent progress on spontaneous scalarization, one can refer to the review Doneva:2022ewd .
Recently, the Einstein-scalar-Gauss-Bonnet (EsGB) theory becomes the most attractive model in which spontaneous scalarization can occur Silva:2017uqg ; Berti:2020kgk ; Herdeiro:2020wei ; Antoniou:2017acq ; Doneva:2017bvd ; Antoniou:2017hxj ; Doneva:2018rou ; Collodel:2019kkx ; Herdeiro:2020wei ; Cunha:2019dwb . In this theory, the scalar field couples to the GB invariant, a higher-order curvature term motivated by high-energy physics Zwiebach:1985uq , which also avoids Ostrogradsky instabilities Ostrogradsky:1850fid . Spontaneous scalarization occurs when the vacuum solution, represented by the Kerr BH, undergoes a tachyonic instability under scalar perturbations as the model parameters exceed a certain threshold Doneva:2020nbb ; Dima:2020yac , leading to the dynamic formation of a stationary scalar field profile near the event horizon East:2021bqk . As a viable model that mimic the Kerr phenomenology, the scalarized BHs in EsGB theory have been studied in different aspects, including the BH shadow Cunha:2019dwb , the dynamical evolution East:2021bqk and the stability analysis Blazquez-Salcedo:2020rhf ; Blazquez-Salcedo:2020caw ; Blazquez-Salcedo:2018jnn ; Silva:2018qhn ; Macedo:2020tbm .
A straightforward approach to test the EsGB theory is measuring the effects of scalarization on GWs of coalescing NSs or BHs. The constraint on EsGB theory with inspiral GW data by considering the absence of scalar dipole radiation has been performed in Wong:2022wni . Nonetheless, the waveform models for the merger and ringdown stages remain underdeveloped. The ringdown waveform can be precisely modeled by the superposition of damped sinusoids with complex eigenfrequencies known as quasinormal modes (QNMs) Kokkotas:1999bd ; Berti:2009kk ; Konoplya:2011qq . The real and imaginary part of each mode, which represents the oscillation frequency and the inverse of the damping time, uniquely determined by the mass and angular momentum of the BH remnant Berti:2005ys .
To perform ringdown tests, the QNM spectrum of rotating scalarized BHs must be determined in advance. The QNMs in EsGB theory have been calculated, but only within the static limit Blazquez-Salcedo:2020rhf ; Blazquez-Salcedo:2020caw . However, calculating QNMs for a general stationary axisymmetric BH remains a recognized challenge. The BH background is no longer classified as Petrov type D; hence, the underlying assumptions of the Teukolsky formalism Teukolsky:1973ha are invalid, and it cannot be directly applied without modifications. In recent years, significant efforts have been devoted to deriving perturbation equations and the calculation of QNMs for rotating BHs within the framework of modified gravity theories, including modified Teukolsky formalism Li:2022pcy ; Hussain:2022ins , full numerical relativity simulations Okounkova:2019zjf ; AresteSalo:2022hua and small-spin approximations Wagle:2021tam ; Pierini:2022eim . In particular, a novel method relying on spectral decomposition has been introduced, through which highly accurate QNMs can be directly obtained by solving the equations governing metric and scalar field perturbations Chung:2023zdq ; Chung:2023wkd ; Blazquez-Salcedo:2023hwg ; Chung:2024ira ; Blazquez-Salcedo:2024oek .
In this paper, we study the QNM spectrum of rotating scalarized BHs in the EsGB theory. Our calculations are based on fully non-perturbative methods, where both the background solutions and perturbation functions are numerically constructed by the spectral decomposition. The corresponding dissipative boundary conditions at the event horizon or the radial infinity are presented in a compact form for the first time. We concentrate on the $l=m=2$ fundamental mode, which is the primary excitation of the ringdown phase for nearly equal-mass binary mergers validated by numerical relativity Buonanno:2006ui ; Berti:2007fi .
@@
00 0 $=$ $\mathcal{E}_{\mu\nu}^{(0)}+\epsilon\mathcal{E}_{\mu\nu}^{(1)}+\mathcal{O}(\epsilon)^{2},$
00 0 $=$ $\mathcal{S}^{(0)}+\epsilon\mathcal{S}^{(1)}+\mathcal{O}(\epsilon)^{2},$ (2)
@@
by following Doneva:2017bvd ; Cunha:2019dwb ; Wong:2022wni . The existence domain Cunha:2019dwb ; Wong:2022wni indicates that BHs with moderate spin ($\chi\approx 0.6$, $\chi$ is the dimensionless spin of BHs) or lower must undergo scalarization if their characteristic scale is comparable to $\lambda$. In contrast, rapidly rotating BHs remain effectively indistinguishable from Kerr BHs predicted by GR. The coupling function with the form of (3) is known as the curvature-induced scalarization. Another model called spin-induced scalarization predicts scalarization exclusively for rapidly rotating BHs, achieved by simply multiplying the coupling function in (3) by $-1$ Herdeiro:2020wei ; Berti:2020kgk .
@@
Figure 1: The migration of the $l=m=2$ fundamental mode for scalarized static BHs on the complex plane with increasing $\lambda/M$. The crosses denote the modes for the gravitational perturbations (black) and the scalar perturbation (green) of the vacuum Schwarzschild BH at $\lambda=0$, respectively. As $\lambda/M$ increases, the tensor mode remains unchanged and the scalar mode varies, as depicted by the green solid line. Scalarized BHs emerge at the threshold $\lambda_{0}/M\approx 1.704$.
@@
Figure 2: The spectrum of the fundamental $l=|m|=2$ QNMs for scalarized Kerr BHs are shown while varying BH spin $\chi$ and fixing the coupling constant $\lambda/M=1.75$. The two gray lines depict the prograde mode ($m=2$) with larger imaginary part and the retrograde mode ($m=-2$) with lesser imaginary part for Kerr BHs.
@@
where $f(r,\theta)$ represents an arbitrary component of ($h_{0},h_{1},H_{0},H_{1},H_{2},K,\phi^{(1)}$), $\Omega_{H}\equiv\frac{W}{r^{2}}|_{r=r_{H}}$ is the angular velocity at the event horizon and $\kappa_{H}\equiv\frac{1}{2r_{H}}e^{F_{0}-F_{1}}|_{r=r_{H}}$ is the surface gravity. The boundary condition at the event horizon can be derived by ensuring the regularity of perturbations there Teukolsky:1973ha ; Xiong:2024 , and can be verified through the perturbation equation of a test scalar field as suggested by Pierini:2021jxd . These boundary conditions are consistent with those of the vacuum Kerr BHs and the static scalarized BHs Blazquez-Salcedo:2020rhf .
@@
with the $\alpha$-derivative matrix $[D^{(\alpha)}]^{(x)}_{ij}\equiv d^{\alpha}C^{(x)}_{i}(x_{j})/dx^{\alpha}$ and $[D^{(\alpha)}]^{(\theta)}_{ij}\equiv d^{\alpha}C^{(\theta)}_{i}(\theta_{j})/d\theta^{\alpha}$ with respect to $x$ and $\theta$, respectively.
@@
Figure 3: The $l=m=2$ modes with different dimensionless coupling constant $\lambda/M$ for rotating scalarized BHs. The values of the fixed BH spin are shown at the top of each column. The axial-led, polar-led and scalar-led modes are depicted by the blue, yellow and green lines, respectively.
@@
For scalarized Kerr BHs, the QNM modes inherit their names from those of the static limit and are named as axial, polar and scalar-led modes Blazquez-Salcedo:2024oek , even though the perturbation components are coupled in general.
@@
Fig.2 illustrates the variation of the fundamental $l=|m|=2$ modes as the spin increases for scalarized BHs with $\lambda=1.75$, including both prograde modes ($m\omega_{R}>0$) and retrograde modes ($m\omega_{R}<0$). It is clear that the isospectrality which holds for the Kerr axial-led and polar-led modes is broken once the BHs are scalarized. We find an interesting behavior for the polar-led mode: within certain spin, the prograde mode can exhibit a smaller imaginary part compared to the retrograde mode. In contrast, the behavior of the two branches for the axial-led modes resembles that of Kerr BHs, where the prograde modes always dominate. Moreover, the real part of prograde mode shows a larger deviation from that of the Kerr BH, compared to the retrograde mode.
@@
We demonstrate the QNMs of rotating scalarized BHs for varying coupling constant $\lambda/M$ and fixing the dimensionless spin $\chi$ in Fig.3. At low spin $\chi=0.2$, the behavior of each family is similar to that in the static BH limit as shown by Fig.1. The polar-led mode dominates only at intermediate $\lambda/M$, while the axial-led mode dominates otherwise. For larger spin $\chi=0.4$, the polar-led mode becomes dominant over larger values of the coupling constant, and the polar-led mode can even maintain dominance for $\chi=0.6$. The scalar-led mode becomes dominant only at either small or large values of the coupling constant, with the range of dominance being relatively narrow.
Discussion — QNMs of rotating BHs are crucial to the construction of ringdown waveform of binary BH mergers. In this paper, we have presented the first full calculation of the QNM spectrum of scalarized Kerr BHs in the EsGB theory, where both the BH solution and perturbations were obtained exactly. We found that the QNM with specific $l,m$ splits into three families: the axial-led, the polar-led and the scalar-led modes. The QNM for each family branch from that of vacuum BHs at the bifurcation point of $\lambda$, where the scalarized BHs are close to the vacuum BHs. We have demonstrated the behavior of $l=m=2$ fundamental modes as the BH spin or the coupling parameter $\lambda$ is varying.
The splitting of QNMs, especially the isospectrality breaking of tensor modes, presents a promising avenue for studying the ringdown phase of BHs. In GR, the $l=m=2$ fundamental mode dominates the ringdown following a merger. Assuming this dominance persists in the EsGB theory, a Bayesian analysis of the $l=m=2$ fundamental mode could uncover new physics near the BH horizon. Although detecting higher-order modes in GW signals remains challenging, this mode alone provides four independent constraints, making it a powerful probe.

### v2

@@ Abstract
Recent studies have shown that rotating black holes can undergo spontaneous scalarization, leading to deviations from general relativity in the strong-field regime. We present the first nonperturbative calculation of the quasinormal modes (QNMs) of scalarized Kerr black holes in Einstein-scalar-Gauss-Bonnet gravity, without assuming small spin or weak coupling. Our results reveal a universal splitting of the fundamental $l=m=2$ mode into axial-led, polar-led, and scalar-led branches, breaking the isospectrality characteristic of Kerr black holes. This splitting offers distinct signatures in the ringdown phase of gravitational wave signals and provides a new avenue to test gravity in the strong-field regime. Our findings open the possibility of probing beyond-GR physics using precision measurements of black hole ringdowns in upcoming gravitational wave observations.
@@
Introduction — The advent of gravitational wave (GW) astronomy has opened a new window into the strong-field regime of gravity, enabling direct tests of general relativity (GR) in previously inaccessible regimes LIGOScientific:2018mvr ; Berti:2015itd . Among GR’s most striking predictions, black holes (BHs) serve as ideal laboratories due to their simplicity and universality Barack:2018yly . In GR, the no-hair theorem asserts that stationary BHs are fully characterized by mass, spin, and electric charge Herdeiro:2015waa . Precise measurements of GW signals from binary mergers now allow tests of this paradigm, including potential deviations arising from additional scalar or other matter fields Isi:2019aib ; Barausse:2020rsu ; Brito:2017wnc ; Hannuksela:2018izj , offering a powerful probe of physics beyond Einstein’s theory.
Spontaneous scalarization is a nonlinear mechanism through which BHs can acquire scalar hair in the nonlinear regime via coupling between a scalar field and a source term, while remaining consistent with weak-field tests Antoniou:2017acq ; Doneva:2017bvd ; Antoniou:2017hxj ; Doneva:2018rou ; Collodel:2019kkx ; Herdeiro:2020wei ; Silva:2017uqg ; Berti:2020kgk ; Herdeiro:2018wub ; Silva:2017uqg ; Cunha:2019dwb . Among various realizations, Einstein-scalar-Gauss-Bonnet (EsGB) gravity has emerged as a particularly compelling framework, where the scalar field couples to the GB term—a high-energy-motivated curvature invariant free from Ostrogradsky instabilities Zwiebach:1985uq ; Ostrogradsky:1850fid . In this theory, scalarization is triggered when Kerr BHs become unstable against scalar perturbations beyond a critical coupling, leading dynamically to the formation of scalarized rotating BHs Dima:2020yac ; Doneva:2020nbb ; East:2021bqk . These solutions closely mimic Kerr geometry but can exhibit observable deviations, especially in their GW emission. Scalarized BHs in EsGB theory have been studied in contexts ranging from shadows Cunha:2019dwb and dynamical formation East:2021bqk to stability and perturbative properties Blazquez-Salcedo:2020rhf ; Silva:2018qhn ; Macedo:2020tbm ; Antoniou:2024gdf . For a recent review, see Doneva:2022ewd .
A natural avenue to test EsGB gravity is through GW observations of coalescing compact binaries, where scalarization can leave distinct imprints. Current constraints focus primarily on the inspiral stage, where the absence of scalar dipole radiation imposes bounds on the coupling parameters Wong:2022wni . However, the merger and ringdown phases remain largely unexplored due to the lack of accurate waveform models in this regime.
The ringdown signal is governed by a superposition of damped sinusoids—quasinormal modes (QNMs)—whose complex frequencies encode the properties of the remnant black hole Kokkotas:1999bd ; Berti:2009kk ; Konoplya:2011qq . To probe EsGB deviations during ringdown, the QNM spectrum of rotating scalarized BH must first be computed. Previous studies have addressed the static case Blazquez-Salcedo:2020rhf ; Blazquez-Salcedo:2020caw , but extending to rotating backgrounds remains challenging: the scalarized Kerr geometry lacks Petrov type D symmetry, invalidating the standard Teukolsky formalism Teukolsky:1973ha . Recent advances have explored alternative methods to compute QNMs in modified gravity, including generalized Teukolsky approaches Li:2022pcy ; Hussain:2022ins , full numerical relativity simulations Okounkova:2019zjf ; AresteSalo:2022hua , and small-spin approximations Wagle:2021tam ; Pierini:2022eim . Notably, a spectral decomposition framework has emerged as a powerful tool to solve coupled perturbation equations for rotating black holes, yielding highly accurate QNMs in a fully nonperturbative manner Chung:2023zdq ; Chung:2023wkd ; Blazquez-Salcedo:2023hwg ; Chung:2024ira ; Blazquez-Salcedo:2024oek .
In this work, we present the first fully nonperturbative calculation of the quasinormal mode (QNM) spectrum for rotating scalarized black holes in EsGB gravity. Utilizing spectral decomposition, we numerically construct both the scalarized background and its linear perturbations, and recast the QNM problem as a generalized eigenvalue system. For the first time, we provide a compact and universal formulation of the dissipative boundary conditions at the event horizon and at spatial infinity. Focusing on the $l=m=2$ fundamental mode—the dominant ringdown excitation in nearly equal-mass binary mergers confirmed by numerical relativity Buonanno:2006ui ; Berti:2007fi —we find that it generically splits into three distinct families: axial-led, polar-led, and scalar-led. This spectral splitting breaks the isospectrality of Kerr BHs in GR and introduces potential observational signatures of BH scalarization. These findings provide a concrete, model-specific prediction of spontaneous scalarization that could be tested in future GW observations. Our work thus opens new possibilities for probing beyond-GR physics through precision BH spectroscopy.
@@
$0$ $=$ $\mathcal{E}_{\mu\nu}^{(0)}+\epsilon\mathcal{E}_{\mu\nu}^{(1)}+\mathcal{O}(\epsilon)^{2},$
$0$ $=$ $\mathcal{S}^{(0)}+\epsilon\mathcal{S}^{(1)}+\mathcal{O}(\epsilon)^{2},$ (2)
@@
by following Doneva:2017bvd ; Cunha:2019dwb ; Wong:2022wni . The existence domain Cunha:2019dwb ; Wong:2022wni indicates that BHs with moderate spin ($\chi\approx 0.6$, $\chi$ is the dimensionless spin of BHs) or lower must undergo scalarization if their characteristic scale is comparable to $\lambda$. In contrast, rapidly rotating BHs remain effectively indistinguishable from Kerr BHs predicted by GR. The coupling function with the form of (3) is known as the curvature-induced scalarization. Another model called spin-induced scalarization predicts scalarization exclusively for rapidly rotating BHs, achieved by simply multiplying the coupling function in (3) by 1-1 Herdeiro:2020wei ; Berti:2020kgk .
@@
Figure 1: The migration of the $l=m=2$ fundamental mode (dashed lines) for scalarized static BHs on the complex plane with increasing $\lambda/M$ described by the direction of arrows. The black (green) crosses denotes the mode for the gravitational perturbation (the scalar perturbation) at $\lambda=0$. As $\lambda/M$ increases, the gravitational mode of the vacuum BH remains unchanged, and the scalar mode varies as depicted by the green solid line. Scalarized BHs branch from the threshold of nonvanishing $\lambda/M\approx 1.704$, and hence the scalar mode for the scalarized branch (the green dashed line) emerges from the green solid line at the point $\omega\approx 0.4661-0.0988i$, instead the green cross (about 0.48360.0968i0.4836-0.0968i). The other dashed lines describe gravitational modes (axial-led modes (blue) and polar-led modes (yellow)) branching from the black cross, where $\lambda/M$ ranges from $1.704$ to $6$. These results exhibit the resemblance of the perturbative response between the vacuum BH and the scalarized BH at the threshold, analogous to their thermodynamic behavior Doneva:2017bvd .
@@
Figure 2: The spectrum of the fundamental $l=|m|=2$ QNMs for scalarized Kerr BHs are depicted while varying BH spin $\chi$ and fixing the coupling constant $\lambda/M=1.75$. The left panel shows the real part of the QNMs with varying spin, while the right panel displays the imaginary part. Solid and hollow markers represent the prograde and retrograde mode of the scalarized BHs, respectively. The different colors of markers corresponding to different modes are listed in the legend within the right panel. For comparison, the grey dashed line show the prograde modes of Kerr BHs with $m=2$, while the retrograde modes with m=2m=-2 are depicted by the grey solid line. From the right panel, the prograde families always dominate the spectrum as the mode with smallest imaginary part, analogous to the spectrum of Kerr BHs (the grey dashed line is always larger than the grey solid line). However, the polar prograde modes (the yellow solid markers) remarkably exhibit a smaller imaginary part than the polar retrograde mode (the purple hollow markers).
@@
where $f(r,\theta)$ represents an arbitrary component of ($h_{0},h_{1},H_{0},H_{1},H_{2},K,\phi^{(1)}$), $\Omega_{H}\equiv\frac{W}{r^{2}}|_{r=r_{H}}$ is the angular velocity at the event horizon and $\kappa_{H}\equiv\frac{1}{2r_{H}}e^{F_{0}-F_{1}}|_{r=r_{H}}$ is the surface gravity. The boundary condition at the event horizon can be derived by requiring the regularity of perturbations in that region Xiong:2024 . Moreover, these conditions have been shown to be universal for all massless perturbations on the background metric given in Eq. (4) Xiong:2024 , and are consistent with those established for vacuum Kerr black holes and scalarized BHs Blazquez-Salcedo:2020rhf ; Pierini:2021jxd ; Chung:2024vaf .
@@
with the $\alpha$-derivative matrix $D^{(\alpha)}]^{(x)}_{ij}\equiv d^{\alpha}C^{(x)}_{i}(x_{j})/dx^{\alpha}$ and $D^{(\alpha)}]^{(\theta)}_{ij}\equiv d^{\alpha}C^{(\theta)}_{i}(\theta_{j})/d\theta^{\alpha}$ with respect to $x$ and $\theta$, respectively.
@@
Figure 3: The $l=m=2$ prograde modes with varying dimensionless coupling constant $\lambda/M$ for rotating scalarized BHs with different spin. The different markers denote the axial modes (blue circles), the polar modes (yellow squares) and the green modes (green diamonds), respectively, as listed in the small frame within the left bottom and held to the remaining panels. The black crosses represent the $l=m=2$ gravitational modes of Kerr BHs with following spin. The three values ($\chi=0.2,0.4,0.6$) of the fixed BH spin are shown at the top of the three column. For each column, the first row displays the real part of QNMs, while the second row shows the imaginary parts. These plots illustrate the emergence of the polar and axial modes for scalarized BHs from the gravitational modes of Kerr BHs at the critical $\lambda/M$. As $\lambda/M$ increases, the polar and axial modes split from each other into two distinct line.
@@
Fig.2 illustrates the variation of the real part (left panel) and imaginary part (right panel) for the fundamental $l=|m|=2$ modes as the spin increases for scalarized BHs with $\lambda=1.75$, including both prograde modes (defined by $m\omega_{R}>0$ and denoted by solid markers) and retrograde modes (defined by $m\omega_{R}<0$ and denoted by hollow markers). For scalarized Kerr BHs, the prograde modes (or the retrograde modes) can further be categorized by three families (the axial, polar and scalar-led modes) according to their main perturbed components (5) Blazquez-Salcedo:2024oek , which their names inherit from those of the static limit Blazquez-Salcedo:2020caw ; Blazquez-Salcedo:2020rhf . In summary, the fundamental modes for given ($l,|m|$) are categorized into six families, with the corresponding markers listed in the legend on the right panel of Fig.2. we also depict the prograde modes (the grey dashed line) of Kerr BHs with $|l|=m=2$ and the retrograde modes (the grey solid line) with |l|=m=2|l|=m=-2 in each panel for comparison. It is clear that the isospectrality, which holds for the Kerr axial-led and polar-led modes, is broken once the BHs are scalarized. We find an interesting behavior for the polar-led modes: for certain spin, the prograde mode (the yellow squares) can exhibit a smaller imaginary part compared to the retrograde mode (purple inverted triangles). In contrast, the behavior of the two branches for the axial-led modes resembles that of Kerr BHs, where the prograde modes always dominate.
@@
We demonstrate the QNMs of rotating scalarized BHs for varying coupling constant $\lambda/M$ and fixing the dimensionless spin $\chi$ in Fig.3. Three fixed values ($\chi=0.2,0.4,0.6$) of the spin are shown at the top of each column for the panels. For each column, the first row presents the real part of the QNMs, while the second row demonstrates the imaginary part. In each panel, the blue circles, yellow squares, and green diamonds represent the axial-led, polar-led, and scalar-led modes, respectively. We depict the $|l|=m=2$ gravitational modes of Kerr black holes by the black cross in each panels, highlighting the splitting between axial-led and polar-led modes for the scalarized BHs. At low spin $\chi=0.2$, the behavior of each family is similar to that in the static BH limit as shown by Fig.1. The polar-led mode dominates only at intermediate $\lambda/M$, while the axial-led mode dominates otherwise. For larger spin $\chi=0.4$, the polar-led mode becomes dominant over larger values of the coupling constant, and the polar-led mode can even maintain dominance for $\chi=0.6$. The scalar-led mode becomes dominant only at either small or large values of $\lambda/M$, with the range of dominance being relatively narrow.
Discussion — We have presented the first nonperturbative computation of the QNM spectrum of rotating scalarized BHs in EsGB gravity. Our approach captures both the nonlinear scalarized background and its full linear perturbations using spectral methods, allowing accurate determination of QNMs beyond previous approximations. A key result is the universal splitting of the fundamental $l=m=2$ mode into three families—axial-led, polar-led, and scalar-led—breaking the isospectrality characteristic of Kerr BHs in GR. This effect, absent in GR, introduces distinct signatures that may be detectable in the ringdown phase of GWs from BH mergers.
Our findings provide concrete theoretical predictions for the observable consequences of BH scalarization. As GW detectors continue to improve in sensitivity, precision spectroscopy of the ringdown phase may offer a novel avenue to probe strong-field gravity and test the limits of Einstein’s theory.